### v2

Abstract
Quantum computing offers a new paradigm for computation, exploiting an exponentially growing Hilbert space for data representation and operation. Computing results are obtained from measurement-based stochastic sampling over non-trivial distributions produced by the quantum computing process with complex correlations from quantum entanglement. Quantum machine learning (QML) emerged as a promising method, potentially surpassing classical counterparts in efficiency and accuracy. While large-scale fault-tolerant quantum (LSFTQ) machines are not yet available, recent works on hybrid quantum machine learning models, compatible with noisy intermediate-scale quantum (NISQ) computers, have hinted at improved generalizability. Such hybrid deep quantum neural networks (hDQNNs) integrate GPU/CPU-based deep neural networks (DNNs) with forward parameterized quantum circuits (PQC) that can be straightforwardly executed on quantum processors. However, backpropagating through forward PQCs remains computationally inefficient if it is simulated on classical hardware, and current quantum hardware constraints impede batched backpropagation even with dedicated PQCs for calculating gradients of forward PQCs, limiting scalability for modern machine-learning tasks. Here, we present a scalable quantum machine learning architecture that overcomes these challenges with efficient backpropagation, enabling scalable hDQNN training with PQCs on physical quantum computers or classical PQC simulators. Applied to reinforcement learning benchmarks, our method highlights hDQNNs’ improved generalizability compared to purely classical models. These findings offer a pathway toward near-term hybrid quantum–classical computing systems for large-scale machine learning and underscore the potential of hybrid quantum-classical architectures in advancing artificial intelligence.
keywords:
Quantum Computing, Machine Learning
1 Introduction
Quantum computing has initially attracted significant attention for its potential to solve classically challenging problems by exploiting data representation and operation in a Hilbert space that grows exponentially with the number of qubits. Among the applications that could benefit most from quantum technologies is machine learning, where increasingly large datasets and embedding dimensions challenge the computational limits of conventional hardware. Quantum machine learning (QML) aims to harness quantum effects, such as superposition and entanglement, to enhance the efficiency and accuracy of learning algorithms beyond classical methods.
Although fault-tolerant quantum machines—capable of running complex circuits without error—are not yet available, the rapid development of noisy intermediate-scale quantum (NISQ) devices has spurred a new class of QML techniques designed to operate under hardware constraints. One motivation came from recent quantum supremacy experiments [1] experimentally demonstrating NISQ quantum devices’ ability to generate distributions that allow sampled output data to represent complex correlations challenging for classical simulations to be efficiently carried out. As generating classical distribution that models the pattern in data is critical for training advanced machine learning models with optimized generalizability [2, 3], NISQ quantum systems’ unique ability to generate nontrivial distribution in sampled output data has the potential to better model the complicated correlations in sophisticated real-world problems. Recent work has shown that such techniques, particularly those integrating classical deep neural networks (DNNs) with parametric quantum circuits (PQCs), can perform on par with classical benchmark reinforcement learning models [4]. This hybrid quantum machine learning is often realized in the form of hybrid deep quantum neural networks (hDQNNs) [4]. Here, a deep neural network (DNN) is coupled to a parameterized quantum circuit (PQC) for forward propagation, seeking to leverage the quantum circuit’s high-dimensional Hilbert space as a feature embedding space while retaining the flexibility and scalability of classical deep learning.
However, a significant barrier to deploying hDQNNs at scale is the computational cost of backpropagating through the PQC. Although specialized methods exist for computing gradients of forward quantum circuits [5], these can be inefficient to scale on classical simulators and often rely on physical quantum hardware that cannot easily support large-batch parallelization. The lack of high-throughput training strategies limits practical performance and applicability to real-world machine-learning tasks.
In this work, we introduce a scalable solution that overcomes the inefficiencies of backpropagation through PQCs. By enabling scalable hybrid training with both quantum circuits and classical DNNs, our architecture delivers efficient backpropagation and supports batched training at scale. We demonstrate the approach on reinforcement learning benchmarks, highlighting the improved generalization of hDQNNs relative to purely classical models, and present a roadmap for near-term quantum–classical integration in large-scale machine learning. This article is organized as follows. We will introduce the QML architecture that enables scalable training of hDQNN with classical DNNs interleaved with PQCs in Sec. 2. Following this, Sec. 3 will extend the scalable hDQNN architecture for reinforced learning tasks. Following this, we show the experimental results demonstrating that hDQNNs could outperform comparable classical methods. In the last Sec. 4, we will discuss novel features observed in experiments and outlooks to train significantly larger models.
2 Neural PQC Tangential Approximation
Modern machine learning methods are established with highly efficient backpropagation-based training. Backpropagation utilizes well-known chain rules for vector calculus to infer the contribution of each learnable parameter, e.g., weights and biases of DNN, in the loss function value (error) by reversely traversing the computation graph of the DNN from the final error. This proven learning method relies on each computing operation node in the DNN’s computation graph to have a closed-form analytical expression for its gradient vector operator that can be explicitly predefined and efficiently calculated (in parallel) in training. The backpropagated errors mapped back from the final error through the chain of gradients on the computation graph are used to update the network parameters such that the updated network could have decreased errors between the network’s predictions and targets in the training dataset. This learning process repeats until the final error reaches below an acceptable value prescribed by the developer or arrives at a steady state.
Figure 1: a, a typical hybrid deep quantum neural network (hDQNN) contains a PreDNN connected to a PostDNN via both direct connections and a quantum layer realized as a parametrized quantum circuit. b, hDQNN is exposed with a batch of training data and the PQC’s input and output vectors are recorded into a Replay Buffer as $\mathbf{q}_{\text{i}}$ and $\mathbf{q}_{\text{o}}$ respectively. The buffered training data is then used to train an ancillary deep neural network (qtDNN) to approximate the behaviors of the PQC in the training dataset. c, the trained qtDNN is then used to replace the PQC in the training process to facilitate the batched backpropagation of errors that are used to update the PreDNN and PostDNN at the same time efficiently. Note that we do not update qtDNN in c.
On the other hand, hDQNNs are typically developed using PQCs to act as functional layers without learnable parameters. They connect data and neural layers (NL) in a model, as seen in Fig. 1a. In these cases, PQCs can be called quantum layers (QL). The classical DNNs before and after the QL are respectively called PreDNN and PostDNN in this work. Classical links directly connecting PreDNN and PostDNN are added. These direct links allow classical communication between the DNNs around a QL. Information forwarded from the PreDNN through the classical links could help the PostDNN properly handle the QL’s measured outputs. For these QLs with a given parametrization structure, the input vector to the QL contain elements that are PQCs’ controlling parameters used in defining the quantum gates to be executed. The QL output vector (bistring)’s elements are the (most probable) measurement results of qubits. Note that qubit measurement results are intrinsically stochastic. This is because a QL maps its input data into a high-dimensional Hilbert space spanned by all possible observable qubit states (bitstrings) as basis vectors. Therefore, for a QL containing $N-$qubits, the embedding space is $2^{N}-$dimensional. Measurements on the embedded data effectively project the data onto one of the basis states with a probability depending on the weight of the basis state in the original state. Consequently, obtaining a QL’s output vector is equivalent to a sampling the underlying probability distribution generated by the PQC defined by its input vector. This fundamentally differentiates QLs from traditional functional layers with deterministic mappings from inputs to outputs. QLs convert inputs into probability distributions of bitstrings and outputs are sampled from them.
This difference makes it challenging to backpropagate errors efficiently through QLs. Various strategies have been developed in past years [5]. Most widely used methods obtain the gradient vector of a PQC with respect to its input parameters by varying each input parameter and using repeated PQC executions to estimate its partial derivative with respect to each varied input. Such methods become quickly intractable as the number of input parameters becomes large. Alternatively, for a given PQC structure, there could be a corresponding PQC for each input variable and its output measurement outcome’s expectation value is an estimator for the original PQC’s partial derivative with respect to the input. Despite requiring many repeated measurements that scale with the size of input and output, methods based on explicit PQC gradient estimation are intrinsically unfeasible to backpropagate errors in batches, prohibiting efficient batched training for DNNs connected to quantum layers’ at scale.
This section presents an hDQNN architecture that is scalable for training while allowing QLs executable on physical quantum computers to be interleaved between classical DNNs. As mentioned, quantum layers are incompatible with scalable batched error backpropagation for training the DNNs in the model. Even though the fundamental physics of quantum layers executed on a quantum computer offer a new route to potentially more generalizable models utilizing entropy related to entanglement induced nontrivial correlation between nodes while, at the same time, hindering training hDQNN efficiently, we propose the quantum tangential space learning (qtDNN) architecture that approximates the local behaviors of a QL with a classical DNN trained together with the hDQNN. This provides scalability through classical batched backpropagation and retains the ability to develop more generalizable machine learning models incorporating QLs running on NISQ computers, overcoming the unique trade-offs.
For each PQC (and its associated derivative estimating PQCs if available) in a given hDQNN model, this architecture adds a parallel DNN to approximate, ideally to the first order, the PQC’s stochastic mapping from input parameters to sampled output bitstring as illustrated in Fig. 1b. In this architecture, the DNN locally approximating the PQC is named a quantum tangent deep neural network (qtDNN). In contrast to conventional principles of designing machine learning models that emphasize avoiding over-fitting the training data, qtDNNs are not negatively impacted by overfitting. Instead, the most resource-efficient is training qtDNN to be an accurate local first-order approximation of the target PQC’s behaviors local to the training data. This is because the qtDNN facilitates batched backpropagation as differentiating and backpropagating through classical DNNs are scalable and established. Meaning that it could be desirable to have a qtDNN designed to memorize the behaviors of its target QL locally. Therefore, a qtDNN bridges backpropagated errors at the output of its corresponding QL to the classical PreDNN. To make training a qtDNN more efficient, we add a qtDNN-buffer to each QL that records the input, $\mathbb{q_{i}}$, and sampled output, $\mathbb{q_{o}}$, of the quantum layer as pairs, $(\mathbb{q_{i}},\mathbb{q_{o}})$. This qtDNN-buffer is then used to train the qtDNN repeatedly in batches. This qtDNN-buffer improves the qtDNN training data efficiency and accelerates qtDNN’s convergence to an accurate local approximation to its target quantum layer. Moreover, as qtDNN aims to approximate the expected bitstring from sampling a target quantum layer, we utilize the binary cross entropy (BCE) loss function to train qtDNN against the qtDNN-buffer data.
The training processes, as seen in Fig. 1a-c, are now explicitly separated into the forward exploration , qtDNN-training, and classical DNN training with qtDNN. To be explicit, we denote the original hDQNN with quantum layers as a function $Q:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ where $n$ and $m$ are the dimensions of model input and output, respectively. We further define the qtDQNN training network approximating $Q$ as $Q_{\text{qt}}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$. $Q_{\text{qt}}$ always shares $Q$’s classical model DNNs and replaces quantum layers with qtDNNs. We further define the training data containing input and target pairs as $(\mathbb{x},\mathbb{y})$ and training $Q$ is tuning its DNNs’ parameters such that $Q(\mathbb{x})\simeq\mathbb{y}$.
In the forward exploration process, as seen in Fig. 1a, input $\mathbb{x}$ from the training data is passed through $Q$ one at a time. In this process, the qtDNN-buffer for each qtDNN is populated with the input and output pairs of the corresponding quantum layer defined above. After this exploration process with sufficient entries in the qtDNN-buffer, qtDNNs are trained with buffered data, in the qtDNN-training process, to approximate the responses of QLs local to the data presented. In the following classical DNN training process with updated qtDNN, training data are sampled into batches. Batched inputs are passed through $Q_{\text{qt}}$ to generate the predictions $\mathbb{y}_{\text{pred}}=Q_{\text{qt}}(\mathbb{x})$. The batched errors of the predictions, $\mathbb{y}_{\text{pred}}$, compared with the expected values, $\mathbb{y}$, in the training data batch, are calculated as scalar loss values using a chosen loss function. In the backward process, the batched scalar errors are backpropagated through the computation graph in reverse, following the chain rule and each node’s Jacobian matrix. In the update process after this batched error backpropagation, the parameters (e.g., weights and biases) of all model DNNs shared by $Q$ and $Q_{\text{qt}}$ are updated using appropriately chosen optimizers. In this parameter updating step, qtDNNs are not updated as they are trained and fixed through training against qtDNN-buffer in the exploration process.
3 Quantum Reinforced Learning
The qtDNN-architecture can be readily applied to solve reinforced learning (RL) problems with hDQNN. This section shows a typical RL framework utilizing hDQNN with qtDNN-architecture as illustrated in Fig. 2. We further implement this framework to solve standard reinforced learning problems and demonstrate excellent performance compared to widely adopted classical baseline models.
Figure 2: a, the typical exploration flow for the Agent to interact with the reinforced learning objective’s target environment and obtain experience data to be stored in the replay buffer. b, the particular modular model architecture used for the Actor Model that allows the intermediate mapping $T$ that maps its input vector $\mathbf{q}_{\text{i}}$ to $\mathbf{q}_{\text{i}}$ to be implemented as 1) PQC,2)a fully connected layer,3) uniform binary number generator, and 4) constant $\mathbf{0}$ output, without changing the rest of the model design. This offers a common testbench to compare learning performance of different representative approaches.
For solving RL with machine learning techniques, we can similarly identify the four processes in training: exploration, forward, backward, and update. RL problems are typically defined by its environment $Env$. An environment is fully characterized by its current state $\mathbb{s}$ and its response to an external action vector $\mathbb{a}$ at this current state. The $Env$ responses to $\mathbb{a}$ conditioned on being in $\mathbb{s}$ are the new state $\mathbb{s}^{\prime}$ of $Env$ caused by $\mathbb{a}$ and a scalar reward value, $r$, for $\mathbb{s}\xrightarrow{\mathbb{a}}\mathbb{s}^{\prime}$.
Utilizing DNNs to solve an RL problem typically requires training two models together as illustated in Fig. 3. One is the Actor Model, $A(\mathbb{s})\rightarrow\mathbb{a}$, which suggests an action vector, $\mathbb{a}$, for a given observed current environment state, $\mathbb{s}$, to maximize the total reward, $R=\Sigma r$, over its future interactions with the $Env$ from current state, $\mathbb{s}$. As the $Env$ is generally not analytically predictable or even deterministic, another useful DNN-based model is the Critic Model, $C(\mathbb{s},\mathbb{a})\rightarrow R$. This model aims to self-consistently approximate the maximal total reward of an action $\mathbb{a}$ for a given environment state $\mathbb{s}$. As the Actor Model drives real actions when the trained RL model is applied to real-world open-loop problems, a more generalizable Actor Model will directly impact the real-world tasks. Therefore, we first start by building the Actor Model based on hDQNN and keep the Critical Model purely classical.
[…]
Figure 3: The batched training process flow that incorporates qtDNN training framework for efficiently updating the classical neural layers of the hDQNN model to achieve desired learning goals.
After the exploration process that generates sufficient data entries in the RB, we could then sample a number ($N_{\text{b}}$) of entries randomly from the RB as a batch. As seen in Fig. 3, Elements corresponding to quantum layers’ $\mathbb{q_{\text{i}}}$ and $\mathbb{q_{\text{o}}}$ in the batched entries are extracted to train and update qtDNNs approximating quantum layers of the Actor Model to the first order within the data.
After updating qtDNNs and obtaining an updated quantum tangential approximated Actor Model $A_{\text{qt}}$ with qtDNNs as batch training compatible approximation for quantum layers, we train and update the classical Critic Model using the standard Temporal Difference (TD) learning method. For training $C$ we define a target network $C_{\text{t}}$, which is a time-delayed copy of $C$. We calculate the estimated total reward using two methods. These two methods should converge on the final values as the training for $C$ converges. Therefore, we train the prediction of $C$ consistently in the iterative training rounds. One method is straightforward, using the batched sampled values for $\mathbb{s}$ and $\mathbb{a}$ to predict total reward $R=C(\mathbb{s},\mathbb{a})$. The other method tries to predict the same value using the fact that total reward could also be calculated as $R_{\text{t}}=r+C_{\text{t}}(\mathbb{s}^{\prime},A_{\text{qt}}(\mathbb{s}^{\prime}))$ if the training for $C\simeq C_{\text{t}}$ converges to accurately predicting the total reward $R\simeq R_{\text{t}}$ and $A\simeq A_{\text{qt}}$ also suggests approximately optimal $\mathbb{a}$ that lead to the largest total future reward from $\mathbb{s}$. Training $C$ is to update its parameters such that it minimizes the loss function characterizing the difference between $R$ and $R_{\text{t}}$. Note that it is only necessary to backpropagate errors through $C$ and update its parameters this way. $C_{\text{t}}$ is updated by copying slightly older parameters of $C$ to keep a time delay between them to avoid oscillatory feedback behaviors causing instability in training. $A$ is only updated through batched backpropagation in the following step.
Now, we have updated $C$ and the qtDNNs of $A_{\text{qt}}$ with the classical DNNs shared with $A$ unchanged. In the last step, we train and update the classical DNNs of $A$ ($A_{\text{qt}}$) using the qtDNNs as an approximation for the quantum layers to the first order in the batched training. We forward pass batched $\mathbb{s}$ into $A_{\text{qt}}$ and $C$ for getting the batched predicted optimal total reward, $R=C(\mathbb{s},A_{\text{qt}}(\mathbb{s}))$. The parameters of only the classical DNNs of $A$ ($A_{\text{qt}}$) are updated to maximize the $R$ using batched backpropagation through $C$ and $A_{\text{qt}}$.
In the rest of this section, we show an experiment implementing qtDNN-based QRL for solving the Humanoid-v4 environment provided by the standard Gymnasium library [6]. We constructed the model in this experiment by modifying Twin Delayed DDPG (TD3) [7, 8]. Allowing for a direct comparison between quantum and classical alternatives based on the same TD3 architecture, we constructed the Actor model as shown in Fig. 2a-b and kept the Critic model purely classical. The classical parts of Actor and Critic models are designed and trained with the parameters in Table 1.
In this experiment, as illustrated in Fig. 2b, the Actor model has a generic block $T:\mathbf{q}_{\text{i}}\rightarrow\mathbf{q}_{\text{o}}$ with input vector $\mathbf{q}_{\text{i}}$ and output vector $\mathbf{q}_{\text{o}}$. Different model constructions are compared by implementing the block with appropriate mapping mechanisms from the input to the output. For this experiment, we implemented the block as 1) PQC, 2) fully-connected linear map (FC), 3) uniform random binary generator (RBG), and 4) constant zeros (0L).
The trained hDQNN-based Actor model outperforms the classical models in terms of test rewards, demonstrating a potential advantage in generalizability. The learning process with key indicators is shown in Fig.4. We noticed that both FC and 0L models underperformed PQC and FC models by a large margin in test rewards with the same exploration noise. This observation is further contrasted with the high rewards achieved in the training loop. This observation highlights entropy’s role in encouraging the model to be trained towards a generalizable global optimal in reinforced learning for complex environments [2, 3]. On the other hand, hDQNN-based model outperforms RBG-based model. This further indicates that learning a structured randomness towards the complex correlated dynamics of a complex high-dimensional environment is desired as learning decreases model entropy [2, 3] for gaining information from interacting with a previously unknown environment. We note that a simple experiment was done for the Pendulum-v1 environment provided by the standard Gymnasium library. This environment is significantly simpler with only a three-dimensional observation space and a one-dimensional action space. The results show that hDQNN-based Actor model trained with qtDNN framework performed optimally with saturating scores on par with alternatives. However, in this simpler environment, we observed that RBG-based model underperformed both quantum and classical alternatives (FC and 0L). This further highlights that even though entropy encourages generalizability in reinforced learning, unstructured randomness that can’t be trained to decrease entropy for the environment to be encoded into the model hinders learning efficiently, especially for simpler and deterministic environments.
Figure 4: The total (test) reward for the trained agents to control the humanoid to walk in Humanoid-v4 environment. The solid lines are rolling window smooth over 10 test rounds. Each test round has 10 testing epochs and each round’s averaged epoch reward is shown as dots in the plot. As the tests are carried out periodically during the training, the horizontal axis records the training epoch after which the round of tests is carried out.
Table 1: hDQNN Reinforced Learning Model Configuration for Gymnasium Humanoid-v4 Problem
DNN Design [^1] DNN designs shared between comparable alternatives Hyperparameters [^2] Hyperparameters shared between comparable alternatives
Components Layers Activation $N$ [^3] The number of qubits in PQC,$M$ [^4] The number of layers in PQC $d_{\text{c-link}}$ [^5] The number of classical links directly between PreDNN and PostDNN $lr$ [^6] Learning rate $\gamma$ [^7] The reward discount parameter $\tau$ [^8] Soft target model update rate
PreDNN [376,256,230] Leaky Relu 10 $3e^{-4}$ $5e^{-3}$
qtDNN [220,2048,10] Leaky Relu $3e^{-4}$
PQC 10,10
PostDNN [20,256,17] Leaky Relu 10 $3e^{-4}$ $5e^{-3}$
Critic [[376,256],[273,256,1]] [^9] The Critic Model has two segments similar to [8] and this represents the configuration of the two segments in forward direction. Leaky Relu $3e^{-4}$ 0.99 $5e^{-3}$
[…]
00footnotetext: Note: Other TD3 parameters such as the exploration action policy noises follow https://huggingface.co/sdpkjc/Humanoid-v4-td3_continuous_action-seed5
4 Discussion
During the experiments presented using hDQNN trained with the qtDNN framework in comparison with alternative classical models, we observed novel behaviors that could provide insights into the fundamental dynamics and physics in learning. These insights could be useful for implementing future improvements and applying the hDQNN-qtDNN framework to major machine learning problems including training and inferences related to the frontier large language models (LLMs).
In the experiments to understand hDQNN’s abilities to learn complex environments’ dynamics efficiently, we implemented a reinforced learning model modularity such that hDQNN’s unique PQC layer could be swapped to alternative classical mappings, which offers a direct way to compare their performances. This work demonstrated hDQNN’s potential generalizability advantage in quantum reinforced learning over classical counterparts. To the best of our knowledge, this is also the first time quantum machine learning has been applied successfully to a complex reinforcement learning task, such as solving the Humanoid-v4 problem, which could potentially be applied in advanced robotic controls.
In this work, we have provided a novel framework for efficiently training hybrid Deep Quantum Neural Networks (hDQNN) for practical, complicated tasks. We further applied this framework to reinforced learning and benchmark its performance in the standard complex Humanoid-v4 environment against comparable classical alternatives. The improved generalizability, shown through achieving the highest test rewards, of hDQNN model trained with qtDNN highlights the importance of engineered entropy in learning towards global optimal with potentially improved training data efficiency. Here, we note that the demonstrated performance of the hDQNN in our experiments for learning classical data and classical environment indicates hDQNN’s ability to model complex correlations and randomness in a complicated unknown environment. In principle, such modeling of complicated classical correlated stochastic behaviors in a classical environment could be learned purely through classical models, and there might indeed be classical models inspired by this work to solve specific reinforced learning tasks with better generalizability and data efficiency. The merit of hDQNNs trained with qtDNN is that hDQNNs have a higher degree of expressivity and could model correlated stochastic processes that are computationally hard for classical models to model [1], while it is straightforward for (noisy) quantum hardware to execute hDQNN in runtime. Training hDQNN with qtDNN framework and inference with hDQNN on (noisy) quantum hardware offer a new route in machine learning and reinforced learning research for addressing complicated real-world problems that are always stochastic, highly correlated, and even challenging to describe with low computing complexities.